in the city, watchingForE continuous inf of in operating rooms at 27. Number of street corner mailbox s Relative 46. A co tc P Survey). Philadelphia 28. Relative humidity levels in oper hospitals. 47. A ber of bananas in a bunch at several flu shot at supermarkets. es of women when they were first ma 48. 31. Weights of the backpacks of first-graders on a school bus the s 85. 32. Namber of students each day who make a with a math tutor at a local college. 33. Duration of marriages in America (in years). de om34. Ages of children in a preschool. 45 For Exercises 35-38, give the boundaries of each val 35. 56 yards. 36. 105.4 miles. 37. 72.6 tons. 38. 9.54 millimeters. ue inal- Section 1-3 For Exercises 39-44, classify each sample as random.S? tematic, stratified, cluster, or other 33. In alarge school district,all teachers from two bul students have less homework to do now thail are interviewed to determine whether they years ether they believe
Answer #1

Variables that are countable (their values can be whole number only, no decimals) are discrete variables and variables that are measurable (their values can be in decimals also) are called continuous variables.

#27 Discrete because Number of street corner mailboxes are countable.

#28 Continuous because relative humidity is measurable.

#29 Discrete because number of bananas are countable.

#30 Continuous because ages are measurable.

#31 Continuous because weights are measurable.

#32 Discrete because number of students are countable.

#33 Continuous because durations are measurable.

#34 Continuous because ages are measurable.

If number has no decimal value, then add and subtract 0.5 from that number to get class boundaries. If number has one digit after decimal value, then add and subtract 0.05 from that number to get class boundaries. If number has two digits after decimal value, then add and subtract 0.005 from that number to get class boundaries, and so on.

#35 The boundaries for 56 yards are:

56-0.5=55.5 and 56+0.5=56.5

#36 The boundaries for 105.4 miles are:

105.4-0.05=105.35 and 105.4+0.5=105.45

#37 The boundaries for 72.6 tons are:

72.6-0.05=72.55 and 72.6+0.05=72.65

#38 The boundaries for 9.54 millimeters are:

9.54-0.005=9.535 and 9.54+0.005=9.545

#39 Cluster Sampling: In cluster sampling, instead of selecting all the subjects from the entire population right off, the researcher takes several steps in gathering his sample population. First, the researcher selects groups or clusters, and then from each cluster, the researcher selects the individual subjects by simple random sampling.

Here all teachers are divided by buildings (clusters) and then randomly select two buildings (clusters) and all teachers from those two selected buildings are interviewed.